A quantity of steam (275g) at 110 degrees C is condensed and the resulting water is frozen into ice at 0 degrees C. How much heat was removed? Solve using kcal. Please solve by showing steps, I'm getting lost somewhere!
Answer #1

A mean calorie is 1?100 of the amount of energy required to warm one gram of air-free water from 0
Answer #2

1. Change temp from 110 to 100.
2. Convert steam into water.
3. Change temp from 100 to 0.
4. Convert water into ice.

1. 275 * .48 * 10 = 1320
2. 275 * 540 = 148500
3. 1 * 275 * 100 = 27500
4. 275 * 80 = 22000

1320 + 148500 + 27500 + 22000 = 199320 cal = 199.32 kcal